OpenStudy (anonymous):

What is the d value of the following arithmetic sequence?27, 41, 55, 69, 83, 97

OpenStudy (helder_edwin):

if it is an arithmetic sequence, then what happens with the difference of two consecutive elements? check that out!

OpenStudy (helder_edwin):

are u waiting for someone to do your homework or something?

OpenStudy (anonymous):

n=n-14 in descending order or n=n+14 in ascending order 27+14= 41 41+14= 55 55+14= 69 69+14= 83 83+14= 97 hence: 97+14=111 or 27-14=13 by the way, in this type of method that Newton invented Calculus.